A balloon containing methane gas has a volume of 2.30 L at 79.0 °C . What volume will the balloon occupy at 158 °C ?
Answer #1

Answer:

Step 1: Explanation:

We know the ideal gas equation, PV=nRT

where, R=universal gas constant,T=temperature ,P=pressure,V=volume,n=moles.

PV=nRT

or, V/T=K [since nR and Pressure all are assume constant]

So, we have V1/T1 = V2/T2.

Step 2: Extract the data from the question

Given, V1 = 2.30 L and  T1= 79 °C = ( 79+273.15 )K = 352.15 K , V2= we need to find and

T2 = 158 °C = ( 158+273.15)K = 431.15 K

Step 3: Calculation

By using above equation

V1/T1 = V2/T2.

or, V2 = V1T2 / T1 = ( 2.30 L × 431.15 K) / 352.15 K= 2.82 L

The new volume is 2.82 L